Find the equation of the line that is parallel to y = 6x – 15 and passes through the origin.
Answer by nerdybill(7384)

y = 6x – 15
is in the "slope-intercept" form
y = mx + b
where
m is slope
b is the y-intercept
.
To be parallel, slopes are identical.
So, our new line has
m = 6
and crosses
(0,0)
.
Plug the above into the "point-slope" form
y - y1 = m(x-x1)
y - 0 = 6(x-0)
y = 6x (this is what they're looking for)